### 03 Mar 2026 v2

# Engineered Simultaneity: The Physical Impossibility of Consolidated Price Discovery Across Spacelike-Separated Exchanges

Abstract
We define engineered simultaneity: the construction of a system that requires temporal comparison of events at spacelike-separated locations, implements this comparison via an implicit simultaneity convention, and represents the result as an objective measurement rather than a conventional choice. We show that the National Best Bid and Offer (NBBO)—the regulatory cornerstone of U.S. equity markets—is an instance of engineered simultaneity. The NBBO requires determining “current” prices across exchanges whose spatial separation places their price events outside each other’s light cones. Special relativity proves that the temporal ordering of such events is frame-dependent: there exist inertial reference frames in which the NBBO differs from the value reported by the Securities Information Processor. The impossibility is not approximate; it is exact and unavoidable within the causal structure of Minkowski spacetime. We formalize the argument using the causal precedence relation, connect it to Lamport’s theorem on distributed ordering, and note that approximately \$5 billion per year in latency arbitrage profits are extracted from the gap between the NBBO’s implicit simultaneity convention and physical reality.
1  Introduction
The U.S. equity market operates across 16 exchanges housed in data centers spanning the New Jersey–Chicago corridor. Regulation NMS, adopted by the Securities and Exchange Commission in 2005 [4], mandates the National Best Bid and Offer (NBBO): a consolidated quotation representing the best available bid and offer prices across all exchanges at any given time. The trade-through rule (Rule 611) requires that trades execute at prices no worse than the NBBO, ensuring investors receive “best execution.”
This regulatory architecture rests on a physical assumption: that “at any given time” has a well-defined meaning across geographically separated exchanges. We show that it does not.
Einstein’s special theory of relativity [1] established that simultaneity is not an absolute relation. For events separated by a spacelike interval—events too far apart in space for light to connect them in the elapsed time—there is no unique answer to the question “which happened first?” Different inertial observers, each applying the laws of physics correctly, assign different temporal orderings to the same pair of events. This is not a measurement limitation. It is a structural feature of Minkowski spacetime [2].
Independently, Lamport [3] proved the equivalent result for distributed systems: without communication, no consistent ordering of events at different nodes can be established. The “happened-before” relation provides the only physically meaningful ordering, and it is a partial order—it leaves spacelike-separated events unordered.
The NBBO requires a total order. Physics provides only a partial order. The gap between these two structures is not a technical deficiency to be engineered away. It is a theorem.
We call the systematic practice of constructing, deploying, and representing such a total order as if it were a physical measurement engineered simultaneity. The purpose of this paper is to define the concept precisely, prove its applicability to the NBBO, and quantify its consequences.
2  The NBBO as a Simultaneity Claim
2.1  Regulatory Definition
Regulation NMS defines the NBBO as the best bid and best offer “currently available” across all “automated quotations” of national securities exchanges [4]. The Securities Information Processor (SIP) receives quotation updates from each exchange and publishes a consolidated NBBO.
The operative word is “currently.” For the NBBO to be well-defined, there must exist a meaningful notion of “current” that applies simultaneously across all exchanges. This requires comparing prices at events—quote updates—occurring at different spatial locations.
2.2  The Consolidation Process
The SIP consolidation proceeds as follows. Each exchange $E_{i}$ ($i=1,\ldots,N$) generates a stream of quote updates. Each update is an event: a price $p_{i}$ produced at spacetime location $(x_{i},t_{i})$. The SIP, located at a fixed position $x_{\mathrm{SIP}}$, receives these updates after propagation delays $\delta_{i}\geq|x_{i}-x_{\mathrm{SIP}}|/c$ and computes:
$\mathrm{NBBO}(t)=\left(\max_{i}\{b_{i}(t)\},\;\min_{i}\{a_{i}(t)\}\right)$ (1)
where $b_{i}(t)$ and $a_{i}(t)$ are the bid and ask prices from exchange $E_{i}$ “at time tt.”
But $b_{i}(t)$ is not the bid at exchange $E_{i}$ at time $t$. It is the most recent bid that has arrived at the SIP by time $t$. The SIP does not observe the market; it observes its own input buffer. The NBBO is a function of arrival times at a single location, not of emission times at multiple locations.
2.3  Formal Definition
Definition 1 (Engineered Simultaneity).
A system exhibits engineered simultaneity when it satisfies three conditions:
  1. (ES1)
    It requires temporal comparison of events at spacelike-separated locations.
  2. (ES2)
    It implements this comparison via an implicit simultaneity convention (typically: order of arrival at a designated consolidation point).
  3. (ES3)
    It represents the result as an objective property of the system rather than as a consequence of the chosen convention.
The NBBO satisfies all three conditions. (ES1): Exchanges are spatially separated and quote events occur at intervals shorter than light-travel time. (ES2): The SIP imposes order by arrival time at its own location. (ES3): Regulation NMS treats the NBBO as the objective “best price” for the purpose of trade-through protection.
3  Causal Structure of the U.S. Exchange Network
3.1  Exchange Locations
U.S. equity exchanges operate from five primary data center clusters. Table 1 lists their locations and pairwise light-time separations.
Cluster Location Exchanges
Mahwah, NJ 41.08N, 74.16W NYSE, NYSE Arca, NYSE American
Carteret, NJ 40.58N, 74.23W Nasdaq, Nasdaq BX, Nasdaq PSX
Secaucus, NJ 40.79N, 74.06W CBOE BZX, BYX, EDGX, EDGA
Weehawken, NJ 40.77N, 74.02W IEX, MEMX, LTSE
Aurora, IL 41.76N, 88.29W CME Group (futures reference)
Table 1: Primary U.S. exchange data center locations.
[…]
Mahwah–Carteret 43 143 215
Mahwah–Secaucus 34 113 170
Carteret–Secaucus 27 90 135
NJ cluster–Aurora 1,180 3,940 5,900
Table 2: Pairwise separations and minimum light-time delays. $\Delta t_{\min}=d/c$ in vacuum; fiber delay assumes $n\approx 1.5$.
[…]
Figure 1: Minkowski diagram of the U.S. exchange network. Events $\alpha$ (NYSE quote) and $\beta$ (CME quote) are spacelike-separated: no light signal from $\alpha$ can reach $\beta$ before $\beta$ occurs. Their temporal ordering is frame-dependent. The SIP, located near the NJ cluster, receives both events and computes an NBBO—but this NBBO reflects arrival order at the SIP, not any frame-independent temporal ordering.
3.2  Minkowski Analysis
Consider two exchanges $E_{A}$ at position $x_{A}$ and $E_{B}$ at position $x_{B}$, with $d=|x_{A}-x_{B}|$. A quote event $\alpha$ at $E_{A}$ occurring at coordinate time $t_{\alpha}$ and a quote event $\beta$ at $E_{B}$ at $t_{\beta}$ are spacelike-separated if and only if:
$|t_{\alpha}-t_{\beta}|<\frac{d}{c}$ (2)
For spacelike-separated events, the Lorentz transformation yields:
$t^{\prime}_{\alpha}-t^{\prime}_{\beta}=\gamma\left[(t_{\alpha}-t_{\beta})-\frac{v}{c^{2}}(x_{\alpha}-x_{\beta})\right]$ (3)
where $v$ is the velocity of the primed frame and $\gamma=(1-v^{2}/c^{2})^{-1/2}$. When events are spacelike-separated, there exist frames in which $t^{\prime}_{\alpha}>t^{\prime}_{\beta}$ and frames in which $t^{\prime}_{\alpha}<t^{\prime}_{\beta}$. The sign of the temporal interval—which event is “first”—depends on the observer.
Theorem 1 (Frame-Dependence of the NBBO).
Let $E_{A}$ and $E_{B}$ be exchanges separated by distance $d>0$. Let $\alpha$ and $\beta$ be quote events at $E_{A}$ and $E_{B}$ respectively, with $|t_{\alpha}-t_{\beta}|<d/c$. Then there exist inertial frames $S$ and $S^{\prime}$ such that the NBBO computed from $\{\alpha,\beta\}$ differs between $S$ and $S^{\prime}$.
Proof.
Suppose $\alpha$ reports bid $b_{A}$ and $\beta$ reports bid $b_{B}>b_{A}$. In frame $S$, where $\alpha$ occurs before $\beta$, the NBBO at any time after both events reflects $b_{B}$. In frame $S^{\prime}$, where $\beta$ occurs before $\alpha$, the sequence of NBBO updates differs: $b_{B}$ enters the NBBO first, followed by no change (since $b_{A}<b_{B}$). More critically, at times between the two events in each frame, the NBBO disagrees: frame $S$ reports $b_{A}$ as best bid while $b_{B}$ has not yet occurred; frame $S^{\prime}$ already reports $b_{B}$.
Since the NBBO is a function of which quote events are deemed “current,” and “current” is frame-dependent for spacelike-separated events, the NBBO is frame-dependent. ∎∎
Corollary 1.
The NBBO reported by the SIP is not a measurement of the market state. It is an artifact of a particular simultaneity convention: the convention that “current” means “received at $x_{\mathrm{SIP}}$ by coordinate time tt.” Different conventions yield different NBBOs, and no convention is privileged by physics.
4  Connection to Distributed Systems Theory
Lamport’s foundational result [3] establishes that in a distributed system, the only physically meaningful ordering of events is the happened-before relation $\rightarrow$, defined by:
  1. 1.
    If $a$ and $b$ are events at the same process and $a$ occurs before $b$, then $a\rightarrow b$.
  2. 2.
    If $a$ is the sending of a message and $b$ is its receipt, then $a\rightarrow b$.
  3. 3.
    The relation is transitive.
Events not related by $\rightarrow$ are concurrent: they have no physically meaningful ordering. Lamport noted explicitly that this structure is isomorphic to the causal structure of special relativity [3]. The happened-before relation is the discrete analogue of the timelike/lightlike ordering in Minkowski spacetime; concurrent events are the analogue of spacelike-separated events.
Proposition 1.
Quote events at spacelike-separated exchanges that have not yet been communicated to any common process are concurrent in Lamport’s sense. The NBBO requires ordering these concurrent events. No such ordering can be derived from the happened-before relation.
The SIP resolves this by fiat: it orders concurrent events by arrival time at a single location. This is a valid synchronization convention but not a unique one. A SIP located elsewhere, or using a different tie-breaking rule, would produce a different NBBO.
Google’s Spanner database [17] confronted the same impossibility. Its solution—TrueTime—represents time as an interval of uncertainty and waits out the uncertainty before declaring a transaction committed. The financial system, by contrast, pretends the uncertainty does not exist.
5  Quantifying the Exploitation
The frame-dependence of the NBBO creates an exploitable asymmetry. Participants who receive exchange data via direct proprietary feeds observe quote events on the order of tens of microseconds after emission [14]; participants relying on the consolidated SIP observe the same events approximately $1{,}128\;\mu\mathrm{s}$ after emission [12]—a ratio exceeding 50:150{:}1.
[…]
This is not a speed advantage in the conventional sense. It is a frame advantage: direct-feed participants effectively inhabit a reference frame closer to the exchange’s local frame, while SIP participants inhabit a frame displaced by the consolidation delay. The two frames assign different temporal orderings to recent quote events, and therefore compute different NBBOs.
[…]
Budish et al. [9] characterize this extraction as socially wasteful: it represents private profit from temporal ambiguity, not compensation for liquidity provision or price discovery.
5.1  The Magnitude Counterargument
Bartlett and McCrary [12] report that the SIP-reported NBBO matches the “true” NBBO 97% of the time, with aggregate gains of only \$14.4 million on \$3.7 trillion in trading volume. Two responses are warranted.
First, Bartlett and McCrary’s “true” NBBO is itself computed using direct-feed data at a single location—another simultaneity convention. They have not measured the gap between convention and reality; they have measured the gap between two conventions.
Second, the argument from magnitude misidentifies the claim. The impossibility result (Theorem 1) does not depend on how often the NBBO is “wrong” or by how much. It depends on whether the NBBO is well-defined. A thermometer that reads correctly 97% of the time but is calibrated to a temperature scale based on phlogiston theory is not approximately right. It is conceptually incoherent. The 3% where it fails is not noise; it is where the incoherence becomes visible, and—per Aquilina et al.—where \$5 billion per year is extracted.
6  Discussion
Engineered simultaneity is a category mistake in the sense of Ryle [18]: it applies a concept (absolute simultaneity) to a domain (spacelike-separated events) where the concept has no meaning. The mistake is not that the SIP computes something—it does. The mistake is that the result is presented as “the best price” rather than as “the best price according to one particular, non-unique simultaneity convention.”
The situation is precisely analogous to the pre-relativistic assumption of absolute time. Newtonian mechanics assumes a universal “now” that all observers share. Special relativity showed this to be a convention, not a fact. Financial regulation has not yet absorbed this lesson.
Three structural observations follow:
First, the impossibility is exact. It does not depend on clock accuracy, synchronization protocols, or future technological improvements. No technology can make spacelike-separated events timelike. The speed of light is not an engineering constraint; it is a law.
Second, the NBBO is not a measurement but a declaration. Timestamps assigned by the SIP do not record when events happened in any frame-independent sense. They record when events were registered under a particular convention. Any system that treats these timestamps as objective measurements of temporal order is committing a category mistake.
Third, the impossibility is not academic. It has measurable economic consequences. Participants with preferential access to the convention’s inputs—direct-feed subscribers physically co-located with exchange matching engines—extract approximately \$5 billion per year from participants who rely on the convention’s output.
We note that Wissner-Gross and Freer [5] identified the relevance of relativistic constraints to financial markets in 2010, Angel [6] argued for regulatory engagement with relativistic constraints, Laughlin, Aguirre, and Grundfest [7] documented the 3 ms latency reduction on the Chicago–New York corridor driven by microwave infrastructure, and Buchanan [8] extended the analysis. The present work differs in formalizing the impossibility as a theorem about the causal structure of the exchange network, rather than as a constraint on arbitrage strategies, and in introducing the concept of engineered simultaneity as a precise descriptor for the mechanism.
7  Conclusion
The NBBO assumes a spacelike hypersurface of simultaneity that special relativity proves cannot exist. We have formalized this as Theorem 1, connected it to Lamport’s impossibility of distributed ordering, and shown that the gap between the NBBO’s implicit simultaneity convention and physical reality is exploited for approximately \$5 billion annually.
The concept of engineered simultaneity—the systematic construction and deployment of systems that assume, implement, and represent impossible temporal comparisons as objective—applies beyond financial markets to any distributed system that treats timestamps as physical measurements of ordering. We submit that recognizing this category mistake is a prerequisite for principled system design.